OpenStudy (lifeisadangerousgame):

Why does 2sin(x) * cos(x) = sin (2x)?

OpenStudy (studygurl14):

It doesn't...?

OpenStudy (lifeisadangerousgame):

It doesn't? I'm supposed to prove that it does:

OpenStudy (marcelie):

its that one identity

OpenStudy (studygurl14):

Oh, sorry. I misread the question. My appologies

OpenStudy (studygurl14):

I read it as = 2sin (x) don't know what I was thinking lol

OpenStudy (lifeisadangerousgame):

I typed it as 2sin (x) by accident, I just edited the original question sorry lol

OpenStudy (studygurl14):

oh, lol. so I'm not going crazy

OpenStudy (lifeisadangerousgame):

nah, you aren't. but I'm supposed to prove that it does, but it doesn't make sense to me,

OpenStudy (kevin):

Why not?

OpenStudy (lifeisadangerousgame):

Because it goes from 2sin(x)cos(x) to just sin (2x), where did the cos(x) go?

OpenStudy (johnweldon1993):

hint* \[\large sin(2x) = sin(x+x)\] What can you do with that?

OpenStudy (steve816):

Replying because I want to know the answer too.

OpenStudy (lifeisadangerousgame):

(x + x) is 2x so sin(2x) = sin(2x)

OpenStudy (johnweldon1993):

Another hint* Remember the addition formula? Being something along the lines of \(\large sin(A+B) = ?\)

OpenStudy (kevin):

As JohnWeldon said, you can change sin(2x) into sin (x + x) Remember formula sin (A + B) = sin A cos B + cos A sin B

OpenStudy (lifeisadangerousgame):

It's still not making sense to me, I've read the addition formula, but none of this was really taught to me by anyone so I don't know any of it

OpenStudy (johnweldon1993):

Don't worry about it! So you know there is an addition formula...which as shown above is: \[\large sin(A+B) = sin(A)cos(B) + sin(B)cos(A)\] So if we apply that here... \[\large sin(x+x) = sin(x)cos(x) + sin(x)cos(x)\] and from there.....

OpenStudy (lifeisadangerousgame):

OH. I see it. sin (x + x) is just sin (2x) which equals sin(x)cos(x) + sin(x)cos(x). Okay, I finally get it. Thank you!

OpenStudy (johnweldon1993):

Exactly! And not a problem!